OpenStudy (anonymous):

Please help me :( 1. A local min of y=x^3+2s^2-1 is? 2.The y-intercept of y=x^3+2x^2-1 is ? 3.The y-intercept of y=x^2+4x-2

OpenStudy (campbell_st):

so in the 1st equation is it \[y = x^3 + 2x^2 - 1\]

Directrix (directrix):

2.The y-intercept of y=x^3+2x^2-1 is ? Substitute 0 for x in the above equation and crank out the value of y for the y-intercept.

Directrix (directrix):

3.The y-intercept of y=x^2+4x-2 Do the same here as for number 2. Suggestion: Check and see if #2 or #3 asked for the x-intercept, okay?

Directrix (directrix):

Is there supposed to be an "s" in this polynomial with the x^3 ? 1. A local min of y = x^3 +2 s^2 - 1